Circular No. 5709

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993E IN KUG 0940+495 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of an

apparent supernova (mag ∼ 19) on a Tech Pan film taken Jan. 27.08 UT with the 0.9-m Schmidt telescope. SN 1993E is located at α = 9^{h}40^{m}26^{s}.26, δ = +49^{o}31′41′′.7 (equinox 1950.0), or 11′′ west and 7′′.8 north of the center of a galaxy known as KUG 0940+495 (Lyon-Meudon extragalactic database). SN 1993E is present at the limiting magnitude of a V plate taken Jan. 17, but not present on the POSS prints. A blue object (mag ∼ 20) has end figures 25^{s}.71, 40′′.6, and a nearby star (mag 15) has end figures 32^{s}.06, 30′45′′.7. A. Maury, J. B. Emond, and Pollas found SN 1993E at mag ∼ 20 on Feb. 14.1 with the same telescope (+ unfiltered CCD).

G272.2–3.2 J. Greiner and R. Egger, Max-Planck-Institut für Extraterrestrische

Physik, Garching, communicate: “G272.2–3.2, a nearly-circular supernova remnant (SNR) newly discovered in the ROSAT All-Sky Survey, is centered at α = 9^{h}06^{m}45^{s}.7, δ = −52^{o}07′03′′ (equinox J2000.0). A thermal plasma model fit reveals a temperature of ∼ 1.6 × 10^{7} K and an interstellar absorption of N_{H} = 2 × 10^{21} cm^{−2} towards G272.2–3.2. This corresponds to ∼ 20 percent of the total galactic column density, suggesting a distance of ∼ 1.5 kpc. Accordingly, the SNR, which has an x-ray radius of ∼ 9′, would be roughly 10^{3} yr old. However, the spectral data also allow a non-thermal origin of the x-ray emission. In particular, G272.2–3.2 is centrally filled and contains an x-ray point source at the geometrical center. Low statistics, however, prevent any time-variability analysis. The x-ray emission surrounding the central 9′ region appears harder than the typical background, indicating that the entire remnant may be larger in size and the central nebula may be synchrotron driven. The pulsar PSR 0905–51 (spindown age 2 × 10^{6} yr) is located near the northeastern rim of G272.2–3.2, suggesting a physical relation between the two objects. The 3σ upper limit for the x-ray emission from PSR 0905–51 is 1.6 × 10^{−13} erg cm^{−2} s^{−1} in the hard band 0.9-2.4 keV and 2 × 10^{−13} erg cm^{−2} s^{−1} in the soft band 0.1-0.4 keV. No radio counterpart of G272.2–3.2 has been found in available radio maps, possibly due to the vicinity of the Vela-SNR complex. On the ESO/SRC R plate, there is a faint nebulosity near the center of G272.2–3.2 and another, more extended one ∼ 6′ west of the center.”

1993 February 15 Daniel W. E. Green